# Extending Graph Burning to Hypergraphs

### v1

@@ 4 Subhypergraphs
Also observe that this example can be extended to provide an infinite family of hypergraphs $H$ with the same property. One may construct $H$ by taking $K_{1,n}$ and adding an edge $e$ containing all $n$ of the vertices of degree one. Then, $b_{L}(H)=1$ and $b(H)=2$. Indeed, by taking the strong subhypergraph of $H$ induced by the vertices in $e$, we obtain a hypergraph that consists of $n$ vertices in an edge, which has lazy burning number $n-1$ and burning number $n$. Thus, the differences in (lazy) burning numbers can be arbitrarily large.
@@
Again, the concept illustrated in this example can be extended to an infinite class of hypergraphs. Let the vertex set of such a hypergraph $H$ be $\{u_{1},u_{2},\ldots,u_{n}\}$, and let the edge set contain $\{u_{1},u_{2}\}$, $\{u_{1},u_{2},u_{3}\}$, $\ldots$ , and $\{u_{1},\ldots,u_{n}\}$. Then, the hypergraph $G$ with vertex set $\{u_{1},\ldots,u_{n-1}\}$ with a single edge containing all of its vertices is a non-induced strong subhypergraph of $H$. Of course, $b_{L}(G)=n-2$ and $b_{L}(H)=1$, so the differences in lazy burning numbers can be arbitrarily large. Now, observe that $b(H)=\left\lceil\frac{n+1}{2}\right\rceil$, since an optimal burning sequence for $H$ is $(u_{1},u_{3},\ldots,u_{n})$ when $n$ is odd, and $(u_{1},u_{3},\ldots,u_{n-1},u_{n})$ when $n$ is even. Since $b(G)=n-1$, the differences in the burning numbers of $G$ and $H$ can also be arbitrarily large.
$H$$e_{1}$$e_{2}$$e_{3}$$u_{6}$$u_{5}$$u_{4}$$u_{3}$$u_{2}$$u_{1}$$u_{7}$$u_{8}$
@@ 5 Discussion and Open Problems
In Section 2 we showed that no analogous bound to the Burning Number Conjecture exists in (lazy) hypergraph burning (i.e. a universal bound on $b_{L}(H)$ or $b(H)$ that is sublinear in terms of $|V(H)|$), even if we only consider hypergraphs that are both uniform and linear. We therefore want to know: what restrictions must we impose on $H$ for such a bound to exist? And what would this bound be? One possibility is that such a bound exists when we insist that $|E(H)|$ is at least $|V(H)|-1$ (which is always the case in a connected graph).
We do not know if either bound in Lemma 3.2 is tight. In Lemma 3.3, we improved the upper bound from Lemma 3.2, assuming that none of the $G_{i}$ are isolated vertices, and this bound is tight. It seems like the bound from Lemma 3.3 also applies to disconnected hypergraphs with up to one isolated vertex, although a proof for this eludes us.
@@
    Define $x(H)$ as the average number of edges a vertex in $H$ belongs to. Does the ratio $\frac{|V(H)|}{x(H)}$ affect $b(H)$ or $b_{L}(H)$? It seems like if $\frac{|V(H)|}{x(H)}$ is large then $b(H)$ is also large. One example of a hypergraph that exhibits this behaviour is the hypergraph consisting of one edge that contains all of its vertices.
@@
    Do there exist stronger bounds or even exact values of $b(H)$ and $b_{L}(H)$ for specific classes of hypergraphs such as projective planes, Kneser hypergraphs, and combinatorial designs? Note that the burning game on Steiner triple systems has been extensively studied in [11].
@@
    Lazy hypergraph burning is in NP. Is it NP-complete?

### v2

@@ 4 Subhypergraphs
Also observe that this example can be extended to provide an infinite family of hypergraphs $H$ with the same property. One may construct $H$ by taking $K_{1,n}$ and adding an edge $e$ containing all $n$ of the vertices of degree one. Then, $b_{L}(H)=1$ and $b(H)=2$. Indeed, by taking the strong subhypergraph of $H$ induced by the vertices in $e$, we obtain a hypergraph that consists of $n$ vertices in an edge, which has lazy burning number n1n-1 and burning number $n$. Thus, the differences in (lazy) burning numbers can be arbitrarily large.
@@
$u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{5}$$e_{1}$$e_{2}$$e_{3}$$e_{4}$$H$
Again, the concept illustrated in this example can be extended to an infinite class of hypergraphs. Let the vertex set of such a hypergraph $H$ be $\{u_{1},u_{2},\ldots,u_{n}\}$, and let the edge set be $\big{\{}\{u_{1},u_{2}\},\{u_{1},u_{2},u_{3}\},\ldots,\{u_{1},\ldots,u_{n}\}\big{\}}$. Then, the hypergraph $G$ with vertex set $\{u_{1},\ldots,u_{n-1}\}$ with a single edge containing all of its vertices is a non-induced strong subhypergraph of $H$. Of course, $b_{L}(G)=n-2$ and $b_{L}(H)=1$, so the differences in lazy burning numbers can be arbitrarily large. Now, observe that $b(H)=\left\lceil\frac{n+1}{2}\right\rceil$, since an optimal burning sequence for $H$ is $(u_{1},u_{3},\ldots,u_{n})$ when $n$ is odd, and $(u_{1},u_{3},\ldots,u_{n-1},u_{n})$ when $n$ is even. Since b(G)=n1b(G)=n-1, the differences in the burning numbers of $G$ and $H$ can also be arbitrarily large.
@@ 5 Discussion and Open Problems
In Section 2 we showed that no analogous bound to the Burning Number Conjecture exists in (lazy) hypergraph burning (i.e. a universal bound on $b_{L}(H)$ or $b(H)$ that is sublinear in terms of $|V(H)|$), even if we only consider hypergraphs that are both uniform and linear. We therefore want to know: what restrictions must we impose on $H$ for such a bound to exist? And what would this bound be? One possibility is that such a bound exists when we insist that $|E(H)|$ is at least |V(H)|1|V(H)|-1 (which is always the case in a connected graph).
We do not know if either bound in Lemma 3.2 is tight. In Lemma 3.3, we improved the upper bound from Lemma 3.2, by assuming that none of the $G_{i}$ are isolated vertices, and this bound is tight. It seems like the bound from Lemma 3.3 also applies to disconnected hypergraphs with up to one isolated vertex, although a proof for this eludes us.
@@
    Do there exist stronger bounds or even exact values of $b(H)$ and $b_{L}(H)$ for specific classes of hypergraphs such as projective planes, Kneser hypergraphs, and combinatorial designs? Note that the burning game on Latin square hypergraphs and Steiner triple systems has been extensively studied in [10] and [13] respectively.
@@
    Lazy hypergraph burning is in NP. Is it NP-complete? (Note that the authors of [11] have since answered this question in the affirmative).
@@
    In [11], the authors find several strong connections between lazy hypergraph burning and a combinatorial process known as zero-forcing. Are there other combinatorial processes that can be linked to (lazy) hypergraph burning?